OpenStudy (anonymous):

Sharon has two solutions available in the lab, one with 6% alcohol and another with 14% alcohol. How much of each should she mix together to obtain 10 gallons of a solution that contains 12% alcohol? I'm having trouble coming up with the equation?

OpenStudy (anonymous):

let x for 6% let y for 14% x+y=10 0.06x+0.14y= 10*0.12

OpenStudy (anonymous):

x+y=10 y=10-x substitution y=10-x 0.06x+0.14(10-x)= 10*0.12 I think u can solve from here

OpenStudy (anonymous):

Thanks for all the help

OpenStudy (anonymous):

0.06x+0.14(10-x)= 10*0.12 0.06x+1.4-0.14x= 1.2 -0.08x = 1.2-1.4 x = -0.2/-0.08 x = 2.5 substitution x= 2.5 find y x+y=10 2.5+y=10 y= 10-2.5 y= 7.5 14% need 7.5 alcohol 6% need 2.5 alcohol you check calculator made sure I'm not calculator wrong